# Permutation-Invariant Spectral Learning via Dyson Diffusion

### v1

@@ 3.1 Dyson Brownian Motion
Dyson (1962) showed that the spectrum of eq. 1 behaves as $n$ positively charged particles in a one-dimensional Coulomb gas. These particles exhibit Brownian motion, but with a pairwise repulsion force proportional to their inverse distance so that their paths do not cross (see Fig. 1c). More precisely, Dyson proved that the spectrum of the entry-wise Ornstein-Uhlenbeck process from eq. 1 follows the SDE given in Theorem 3.1. Without loss of generality, we restrict the domain to the Weyl Chamber $C_{n}:=\left\{\lambda\in\mathbb{R}^{n}:\lambda_{1}>\ldots>\lambda_{n}\right\}.$
Theorem 3.1 (Eigenvalue SDE, Dyson (1962)).
Denote by $\lambda(t)=(\lambda_{1}(t),\ldots,\lambda_{n}(t))$ the ordered spectrum of the matrix-valued Ornstein-Uhlenbeck process $M(t)$ of eq. 1. Then assuming that the initial matrix $M(0)$ has simple spectrum, $\lambda(t)$ satisfies for all $1\leq k\leq n$ the stochastic differential equation
@@
for $W_{1},\ldots,W_{n}$ independent standard Brownian motions. Moreover the unique stationary distribution of (Dyson-BM) has density
for $\lambda\in C_{n}$ and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ corresponds to a probability measure.
For completeness, a full proof of Theorem 3.1 is given in Appendix B, where we generalize a well-known proof to arbitrary coefficients $\alpha,\beta$. We note that the assumption of Theorem 3.1 that $M_{0}$ has simple spectrum is minor. Indeed, generic random graphs or matrices have simple spectra, and in the case of eigenvalues with higher multiplicity, we can perturb the spectrum to be simple.
From Dyson-BM we see that the eigenvalues perform a Brownian motion in a confining potential with a repulsion force: once a pair of eigenvalues $\lambda_{k},\lambda_{l}$ comes too close, they become repelled with a force $\alpha/(\lambda_{k}-\lambda_{l})$ inversely proportional to their separation. A remarkable property of Theorem 3.1 is that the evolution of the spectrum is decoupled from all other information about the matrix: the spectral SDE (Theorem 3.1) is independent of the eigenvectors. This is the key analytical insight that motivates our approach.
Furthermore, conditioned on the eigenvalues, the remaining information captured in form of the eigenvectors can be deduced as we show in Theorem 3.2. This generalizes a statement of Allez et al. (2014), and we give a proof in Appendix C.
Theorem 3.2 (Eigenvector SDE).
Denote by $(v_{1}(t),\ldots,v_{n}(t))$ the orthonormal eigenvectors associated to the eigenvalues of Theorem 3.1. Assuming that the initial matrix $M(0)$ has simple spectrum, $v_{k}(t)$ satisfies for $k\in[n]$ the stochastic differential equation
@@
3.2 From the Dyson SDE to a Diffusion Model
@@
Despite its advantages, constructing a diffusion model based on Dyson-BM poses several challenges, e.g. dealing with a singular drift, non-Gaussian conditional density, etc. (see Appendix J for details). As described below, with DyDM we overcome these obstacles and design an efficient diffusion model for the spectrum, which can distinguish between spectra of graphs that GNNs are blind to (Fig. 2) and which does not require ad hoc data augmentation.
The time-reversal of the Dyson-BM in the sense of Anderson (1982) reads
@@
$\mathrm{d}\bar{\lambda}_{k}(t)=\left[\alpha\sum_{\ell\neq k}\frac{1}{\bar{\lambda}_{k}(t)-\bar{\lambda}_{\ell}(t)}-\beta\bar{\lambda}_{k}(t)-2\alpha[s(\bar{\lambda}(t),t)]_{k}\right]\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}\overline{W}_{k}(t),$ (5)
@@
where we aim to learn the score $s(\lambda,t)\coloneq\gradient_{\lambda}\log p_{t}(\lambda)$. Because the coefficients in Dyson-BM are non-Lipschitz, the applicability of Anderson (1982) is not immediate. Accordingly, in Appendix D we sketch a proof of existence and uniqueness of a strong solution and verify that Anderson’s time reversal applies.
@@
Making the loss tractable. Learning the loss function $s(\lambda,t)$ as in eq. 3 is not feasible for the Dyson-BM, since a closed form of the conditional distribution $p_{t\mid 0}$ is not known in contrast to the OU process. To overcome this, we follow a derivation in the style of De Bortoli et al. (2022) to obtain the loss function – up to constants in $\theta$ – for any $h\in\mathbb{R}^{+}$
@@
$\tilde{L}(\theta)=\operatorname{\mathbb{E}}_{t\sim\mathrm{Unif}[0,T],\lambda_{t}\sim p_{t},\lambda_{t+h}\sim p_{t+h\mid t}(\cdot|\lambda_{t})}\left[\norm{s_{\theta}(\lambda_{t+h},t+h)-\gradient_{\lambda_{t+h}}\log p_{t+h\mid t}(\lambda_{t+h}|\lambda_{t})}^{2}_{2}\right],$ (6)
where we will approximate the intractable $p_{t+h\mid t}$ with the Gaussian transition of the Euler-Maruyama approximation (see Appendix A) for details.
@@
Handling singularities. Numerical solutions of Dyson-BM with a fixed step size are not practical, since Dyson-BM is singular at the boundary of the Weyl Chamber. A fixed step size leads to inaccuracies at the boundaries and may overshoot the singularity, leaving the Weyl Chamber. To overcome this, we implement an adaptive step-size algorithm which conditions on an event of probability $1$ (non-crossing) and hence does not change marginal densities. [^4] Note that the adaptive step-size is only used in the forward simulation, whereas the objective is evaluated always on a fixed grid to ensure correctness. The step-size controller is described in Algorithm 2 and in Section F.1 in detail.
@@
Schedule. Dyson’s conjecture states that for $\alpha=\frac{1}{n}$ and $\beta=\frac{1}{2}$ Dyson-BM converges to the invariant distribution in time $\Theta(1)$, while the majority of eigenvalues mixes locally already in time $\Theta(1/n)$ (Yang, 2022). Hence, the fine-grained structure will be mixed in the time interval $\left(0,\Theta(1/n)\right)$. We show in Appendix E through time change that this conjecture can be applied to any choice of coefficients $\alpha,\beta$ as. It is thus sensible to choose an exponential schedule. We specify the particular choices in Section I.1.
@@
1:Input: spectral samples $\lambda^{(1)},\ldots,\lambda^{(N)}\in\mathbb{R}^{n}$, schedule $\mathcal{T}=\{t_{j}\}$
@@
7:   $\delta t\leftarrow\text{ForwardStepsizeController}(\lambda(t),u)$ $\triangleright$ Conditions on non-intersection, see Appendix F.1.
@@
11:   $\delta t\leftarrow\min\{\delta_{t},t_{\text{fix}}-t\}$ where $t_{\text{fix}}=\min\{t_{j}\in\mathcal{T}:t_{j}>t\}$
12:   $\lambda(t+\delta t)\leftarrow$ Euler-Maruyama step of Dyson-BM with step size $\delta t$ and noise $u$
13:   $t\leftarrow t+\delta t$
14:  end while
15:end for
@@
16:Update $s_{\theta}$ using loss $\tilde{L}(\theta)$ along paths $\lambda^{(1)}(t),\ldots,\lambda^{(N)}(t)$ on schedule $\mathcal{T}$ using eq. 18
17:Output: $s_{\theta}$.
@@
Figure 3: Dyson Diffusion Model (training): The Dyson-BM is evolved forward in time with an adaptive step size ensuring that the paths remain in the Weyl Chamber. The step-size controller conditions on the probability 11 event of non-crossing as detailed in Section F.1 and Algorithm 2.
Equilibrium shooting mechanism. During inference, we require access to the learned score. In the forward diffusion, solely numerical errors due to the time discretisation may lead to crossings of singularities. On the one hand, when going backwards in time, due to inconsistencies in the learned score, the repulsion in eq. 5 might be too weak and the sample path may leave the Weyl Chamber for any sensible step size. Since the score is not defined outside the Weyl Chamber, this is problematic. On the other hand, the step size obtained by conditioning on not leaving the Weyl Chamber in this ill-trained point would be so small that the numerical solver would get stuck. To overcome this, we incorporate a shooting mechanism: If the repulsion force is too weak to prevent crossing of the singularity, resulting in a microscopic step size upon conditioning to remain in the Weyl Chamber, we repel with the invariant-state drift, i.e., we replace the learned score with the score in the invariant state (see Appendix G). This mechanism ensures that we stay in the Weyl Chamber, while minimizing its impact on the distribution. With well-tuned parameters, the shooting gets rarely triggered (less than $0.5\%$ of steps) but is essential, as already a single event would cause getting stuck (upon conditioning) or leaving the Weyl Chamber.
@@
Sampling from Invariant Distribution To sample from the invariant distribution of $\lambda(t)$ given by eq. 4 we exploit the connection between $\lambda(t)$ and $M(t)$: We first sample from the Gaussian invariant distribution of $M(t)$ and then perform an eigendecomposition.
Comparison to direct simulation. One may wonder why we do not sample from the matrix-valued OU process at any time $t$ directly, then perform an eigendecomposition to get $\lambda(t)$, determine $\lambda(t+dt)$, and learn the score network from the increment using eq. 6. This way, learning would be simulation-free and the Dyson SDE would only be needed for the derivation of the loss function as well as the backwards dynamics. The reason for not pursuing this is efficiency and precision. We found that the above way takes $150$ times longer for graphs of size $n=10$, since (accurate) eigendecomposition is computationally costly. Doing so at every step explodes the costs. Moreover, note that we train on the sample of increments, so that from a simulation until time $t$, we learn on the entire generated path. Through our efficient implementation of the SDE, the learning process has running time (up to smaller time steps performed by the adaptive step size controller) on the order of a simulation-free diffusion model.
@@ 4 Experiments
We empirically evaluate DyDM against several state-of-the-art graph diffusion models. First, we compare them on a simple bimodal distribution between two WL-equivalent graphs, illustrating the struggle of purely GNN-based and graph-transformer-based methods (Section 4.1). Next, we carry out a comparison on a standard graph benchmark datasets (Section 4.1) and demonstrate scalability on a larger dataset ($15^{\prime}000$ graphs).
Table 1: Statistical distances of DyDM compared to standard models: DyDM learns the spectrum better in both the $n$-dimensional mean and the marginal Wasserstein sense, without requiring ad hoc data augmentation. Results are rounded to two decimal places, and results ()(*) are equal until the fourth decimal place, and ()(**) until the third decimal place. Exact numbers are provided in Section N.6.
Dataset
WL-Bimodal
Community Small Brain
@@
DyDM (ours) $\bm{0.02}$ $\bm{0.01^{*}}$ $\bm{0.07}$ $\bm{0.02}$ $\bm{0.05}$ $\bm{0.03^{**}}$
EDP-GNN 0.13 0.08 0.42 0.14 - -
GDSS 0.23 0.13 0.42 0.14 - -
ConGress 0.38 0.16 0.27 0.11 - -
DiGress (no trick) 1.06 0.29 2.51 0.45 - -
DiGress (trick) 0.03 $\bm{0.01^{*}}$ 0.09 0.03 0.12 $\bm{0.03^{**}}$

### v2

@@ 3.1 Dyson Brownian Motion
Dyson [24] showed that the spectrum of (1) behaves as $n$ positively charged particles in a one-dimensional Coulomb gas. These particles exhibit Brownian motion, but with a pairwise repulsion force proportional to their inverse distance so that their paths do not cross (see Fig. 1c). More precisely, Dyson proved that the spectrum of the entry-wise Ornstein–Uhlenbeck process from (1) follows the SDE given in Theorem˜2. Without loss of generality, we restrict the domain to the Weyl Chamber $C_{n}:=\left\{\lambda\in\mathbb{R}^{n}:\lambda_{1}>\ldots>\lambda_{n}\right\}.$
Theorem 2 (Eigenvalue SDE [24]).
Denote by $\lambda(t)=(\lambda_{1}(t),\ldots,\lambda_{n}(t))$ the ordered spectrum of the matrix-valued Ornstein-Uhlenbeck process $M(t)$ of (1). Then assuming that the initial matrix $M(0)$ has simple spectrum, $\lambda(t)$ satisfies for all $1\leq k\leq n$ the stochastic differential equation
@@
for $W_{1},\ldots,W_{n}$ independent standard Brownian motions. Moreover, the unique stationary distribution of (Dyson-BM) has density $p_{\mathrm{inv}}(\lambda)=\exp(-U(\lambda))/Z$ for $U(\lambda)=\frac{\beta}{2\alpha}\sum_{k}\lambda_{k}^{2}-\sum_{k<\ell}\ln|\lambda_{k}-\lambda_{\ell}|$, $\lambda\in C_{n}$, and $Z$ a normalizing constant so that $p_{\mathrm{inv}}$ corresponds to a probability measure.
For completeness, a full proof of Theorem˜2 is given in App. B, where we generalize a well-known proof to arbitrary coefficients $\alpha,\beta$. We note that the assumption of Theorem˜2 that $M_{0}$ has simple spectrum is minor. Indeed, generic random graphs or matrices have simple spectra, and in the case of eigenvalues with higher multiplicity, we can perturb the spectrum to be simple. From (Dyson-BM) we see that the eigenvalues perform a Brownian motion in a confining potential with a repulsion force: once a pair of eigenvalues $\lambda_{k},\lambda_{l}$ comes too close, they become repelled with a force $\alpha/(\lambda_{k}-\lambda_{l})$ inversely proportional to their separation. A remarkable property of Theorem˜2 is that the evolution of the spectrum is decoupled from all other information about the matrix: the spectral SDE (Theorem˜2) is independent of the eigenvectors. This is the key analytical insight that motivates our approach. Furthermore, conditioned on the eigenvalues, the remaining information captured in form of the eigenvectors can be deduced as we show in Theorem˜3. This generalizes a statement of Allez et al. [25], and we give a proof in App. C.
Theorem 3 (Eigenvector SDE).
Denote by $(v_{1}(t),\ldots,v_{n}(t))$ the orthonormal eigenvectors associated to the eigenvalues of Theorem˜2. Assuming that the initial matrix $M(0)$ has simple spectrum, $v_{k}(t)$ satisfies for $k\in[n]$ the stochastic differential equation
$\mathrm{d}v_{k}(t)=-\frac{\alpha}{2}\sum_{\ell\neq k}\frac{v_{k}(t)}{(\lambda_{k}(t)-\lambda_{\ell}(t))^{2}}\mathrm{d}t+\sqrt{\alpha}\sum_{\ell\neq k}\frac{v_{\ell}(t)}{\lambda_{k}(t)-\lambda_{\ell}(t)}\mathrm{d}w_{\ell k}(t)$ (Eigenvector-SDE)
@@
3.2 From Dyson SDE to Spectral Diffusion Model
@@
The analytical expression in Theorem˜2 allows learning the spectrum of a matrix-diffusion following (1). Despite its advantages, constructing a diffusion model based on (Dyson-BM) poses several challenges, e.g. dealing with a singular drift, non-Gaussian conditional density, etc. (see App. L for details). As described below, with DyDM we overcome these obstacles and design an efficient diffusion model for the spectrum, which can distinguish between spectra of graphs that GNNs are blind to (Fig. 2) and which does not require ad hoc data augmentation.
The time-reversal of the (Dyson-BM) in the sense of [20] reads
@@
$\mathrm{d}\bar{\lambda}_{k}(t)\!=\!$ $\Bigg[\!\alpha\!\sum_{\ell\neq k}\!\frac{1}{\bar{\lambda}_{k}(t)\!\!-\!\!\bar{\lambda}_{\ell}(t)}\!-\!\beta\bar{\lambda}_{k}(t)\!-\!2\alpha[s(\bar{\lambda}(t),t)]_{k}\!\Bigg]\!\mathrm{d}t+\sqrt{2\alpha}\mathrm{d}\overline{W}_{k}(t),$ (4)
@@
where we aim to learn the score $s(\lambda,t)\coloneq\gradient_{\lambda}\ln p_{t}(\lambda)$. As the coefficients in (Dyson-BM) are non-Lipschitz, the applicability of [20] is not immediate. Accordingly, in App. D we present a proof of existence and uniqueness of a strong solution and verify that Anderson’s time reversal applies.
@@
Making the loss tractable. Learning the loss function $s(\lambda,t)$ as in (3) is not feasible for the (Dyson-BM), since the conditional distribution $p_{t\mid 0}$ is not available in closed form in contrast to the OU process. To overcome this, we follow a derivation in the style of [26] to obtain the loss function – up to constants in $\theta$ – for any $h\in\mathbb{R}^{+}$
@@
$\tilde{L}({\mkern-1.0mu}\theta{\mkern-1.0mu})\!\!=\!\!\mathbb{E}\!\!\left[\!\norm{{\mkern-1.0mu}s_{\theta}(\lambda_{t{\mkern-1.0mu}+{\mkern-1.0mu}h},\!t\!+\!h)\!\!-\!\!\gradient_{\lambda_{t{\mkern-1.0mu}+{\mkern-1.0mu}h}}\!\!\ln p_{t{\mkern-1.0mu}+{\mkern-1.0mu}h|t}(\lambda_{t{\mkern-1.0mu}+{\mkern-1.0mu}h}|\lambda_{t}){\mkern-1.0mu}}^{2}_{2}\!\right]\!,\!$ (5)
the expectation being over $t\sim\mathcal{U}[0,T],\lambda_{t}\sim p_{t},\lambda_{t+h}\sim p_{t+h\mid t}(\cdot|\lambda_{t})$. We approximate $p_{t+h\mid t}$ with the Gaussian transition of the Euler–Maruyama step, which is exact for $h\!\to\!0$ (see App. A).
@@
Handling singularities. Numerical solutions of (Dyson-BM) with a fixed step size are not practical, since (Dyson-BM) is singular at the boundary of the Weyl Chamber. A fixed step size leads to inaccuracies at the boundaries and may overshoot the singularity, leaving the Weyl Chamber. To overcome this, we implement an adaptive step-size algorithm which conditions on an event of probability $1$ (non-crossing) and hence does not change marginal densities. [^4] The adaptive step-size is only used in the forward simulation. The objective is evaluated on a fixed grid for correctness. The step-size controller is described in Algorithm 2 and in App. F.1 in detail.
@@
Schedule. Dyson’s conjecture states that for $\alpha\!=\!\frac{1}{n}$ and $\beta\!=\!\frac{1}{2}$ (Dyson-BM) converges to the invariant distribution in time $\Theta(1)$, while the majority of eigenvalues mixes locally already in time $\Theta(1/n)$ [27]. Hence, the fine-grained structure will be mixed in the time interval $\left(0,\Theta(1/n)\right)$. We show in App. E through time change that this conjecture can be applied to any choice of $\alpha,\beta$. It is thus sensible to choose an exponential schedule. The particular choices are specified in App. K.1.
@@
1:spectral samples $\lambda^{(1)}\!\!\!\!\!,\mkern-1.0mu\ldots\mkern-1.0mu,\mkern-1.0mu\lambda^{(N)}\!\in\!\mathbb{R}^{n}\!\!\!,$ schedule $\mathcal{T}\!\!=\!\!\{\mkern-1.0mut_{j}\mkern-1.0mu\}$
@@
7:   $\delta t\leftarrow\text{ForwardStepsizeController}(\lambda(t),u)$ $\triangleright$ Conditions on non-intersection, see F.1.
@@
11:   $t_{\text{fix}}=\min\{t_{j}\in\mathcal{T}:t_{j}>t\}$
12:   $\delta t\leftarrow\min\{\delta_{t},t_{\text{fix}}-t\}$
[…]
14:   $t\leftarrow t+\delta t$
15:  end while
16:end for
@@
17:Update $s_{\theta}$ using loss $\tilde{L}(\theta)$ along paths $\lambda^{(1)}(t),\ldots,\lambda^{(N)}(t)$ on schedule $\mathcal{T}$ using Eq. (17)
[…]
@@
Figure 3: Dyson Diffusion Model (training): The (Dyson-BM) is evolved forward in time with an adaptive step size ensuring that the paths remain in the Weyl Chamber. The step-size controller conditions on the probability 11 event of non-crossing as detailed in App. F.1 and Algorithm˜2.
Equilibrium shooting mechanism. In the backward dynamics, imperfect learning may erroneously lead to crossing of the Weyl Chamber. To overcome this, we introduce a repulsion force deduced from the equilibrium distribution, acting close to the boundary. For details, see App. G.
@@
Sampling from Invariant Distribution. To sample from the invariant distribution of $\lambda(t)$ given by $p_{\mathrm{inv}}$ we exploit the connection between $\lambda(t)$ and $M(t)$: We first sample from the Gaussian invariant distribution of $M(t)$ and then perform an eigendecomposition.
Comparison to direct simulation. There are clear efficiency and precision reasons for why we do not sample from (1) in combination with eigendecompositions, see App. G.1. Moreover, our SDE implementation renders the training runtime on the order of a simulation-free diffusion model.
3.3 Eigenvector Time Reversal and Numerical Scheme
While DyDM focuses on learning spectra, a defining feature is that the eigenvector information remains accessible. We now present the time reversal and numerical scheme for the (Eigenvector-SDE), requiring novel methods that go beyond existing work on diffusions on the Lie group [26, 28].
From (Eigenvector-SDE), we derive in App. S that $X_{t}=(v_{1}(t),\dots,v_{n}(t))$ follows $X_{t+\mathrm{d}t}=X_{t}\exp(\mathrm{d}Z_{t}(\lambda))$ where $\mathrm{d}Z_{t}(\lambda)=\sqrt{\alpha}\sum_{\ell<k}\frac{E_{\ell,k}\mathrm{d}W_{\ell,k}}{\lambda_{k}(t)-\lambda_{\ell}(t)}$ with $E_{\ell,k}$ being $1$ at the $(\ell,k)$-entry, 1-1 at $(k,\ell)$, and 0 otherwise, and i.i.d. Wiener increments $\mathrm{d}W_{\ell,k}$. Note that $\mathrm{d}Z_{t}(\lambda)^{T}=-\mathrm{d}Z_{t}(\lambda)$ is in the Lie algebra $\mathfrak{o}(n)$ of the Lie group $\mathrm{O}(n)$, and hence the matrix exponential implies that $X_{t+\mathrm{d}t}\in\mathrm{O}(n)$.
The time-reversed diffusion $\bar{X}_{t}$ on the Lie group follows $\bar{X}_{t+\mathrm{d}t}=\bar{X}_{t}\exp(\mathrm{d}\bar{Z}_{t}(\lambda,\bar{X}_{t}))$ with $\mathrm{d}\bar{Z}_{t}(\lambda,\bar{X}_{t})=\alpha\nabla\ln p_{t}(\bar{\lambda},\bar{X}_{t})dt+\sqrt{\alpha}\sum_{\ell<k}\frac{E_{(\ell,k)}}{\lambda_{k}(t)-\lambda_{\ell}(t)}\mathrm{d}\bar{W}^{(\ell,k)}_{t}\in\mathfrak{o}(n)$ as we show in App. S.3.
Phrasing the forward- and backward-dynamics, including the score, via Lie algebra increments $\mathrm{d}Z_{t},\mathrm{d}\bar{Z}_{t}$ offers multiple key advantages. First, orthogonality and normalization of $X(t),\bar{X}(t)$ during simulation follow directly. Crucially, the number of degrees of freedom is n(n1)/2n(n-1)/2 instead of the naive $n^{2}$, as is apparent from the summation over $\ell<k$. Strikingly, it enables to implement a permutation-invariant learning architecture. While permutation equivariance of the columns of $X_{t}$ is already resolved, since they correspond to uniquely ordered eigenvalues, permutation equivariance of the rows of $X_{t}$ is covered by the following Lemma. This represents two levels of simplification, since permutation equivariance of both, rows and columns, is simplified to fixed columns and permutation-invariance in the rows, as we prove in App. S.5.
Lemma 4 (Permutation invariance of score in Lie algebra).
For any permutation matrix $P$ we have $\nabla\ln p_{t}(\bar{\lambda},P\bar{X}_{t})=\nabla\ln p_{t}(\bar{\lambda},\bar{X}_{t})\in\mathfrak{o}(n)$ for all $t\geq 0$, $\bar{\lambda}(t)\in{\rm Weyl\ Chamber}$ and $\bar{X}_{t}\in O(n)$.
Finally, we exploit the symmetry under sign flips of eigenvectors that cannot influence the dynamics. This implies that the score obeys $\nabla\ln p_{t}(\bar{\lambda},\bar{X}_{t}F)=F\nabla\ln p_{t}(\bar{\lambda},\bar{X}_{t})F$ for any diagonal matrix $F$ with entries $\pm 1$, thus a reduction of the learning target of order $2^{d}$, as we prove in App. S.6.
@@ 4 Experiments
We empirically evaluate DyDM against several state-of-the-art graph diffusion models for learning graph spectra. First, we compare them on a simple bimodal distribution between two WL-equivalent graphs, illustrating the struggle of GNN-based methods (Section˜4.1). Next, we carry out a comparison on a standard graph benchmark dataset and demonstrate scalability on a larger dataset (15,00015{,}000 graphs).
Table 1: Statistical distances of DyDM compared to established models: DyDM learns the spectrum better in both the $n$-dimensional mean and the marginal Wasserstein sense, without requiring ad hoc data augmentation. Results are rounded to two decimal places, and results ()(*) are equal until the fourth decimal, and $(\dagger)$ until the third decimal (see App. Q.6).
Dataset WL-Bimodal Comm. Small Brain
@@
DyDM (ours) $\bm{0.02}$ $\bm{0.01^{*}}$ $\bm{0.07}$ $\bm{0.02}$ $\bm{0.05}$ $\bm{0.03^{\dagger}}$
EDP-GNN 0.13 0.08 0.42 0.14 0.07 $0.03^{\dagger}$
GDSS 0.24 0.13 0.38 0.14 0.33 0.12
ConGress 0.38 0.16 0.27 0.11 0.13 $0.03^{\dagger}$
DiGress (no trick) 1.06 0.29 2.51 0.45 0.57 0.17
DiGress (trick) 0.03 $\bm{0.01^{*}}$ 0.09 0.03 0.12 $\bm{0.03^{\dagger}}$
We further demonstrate in App. S the accessibility of the eigenvectors by learning the Lie group SDE’s time reversal of Section 3.3 on $SO(3)$ and with a numerical adjustment we learn a Lie group diffusion for the WL-Bimodal case.